table
cola root beer sprite
under 21 years of age= 40 25 20
between 21 and 40= 35 20 30
over 40= 20 30 35
If one of the 255 subjecta is randomly selected , find the probability that the person is over 40 years of age (fraction form) Please let me know formula used.. I want to learn :)

Answer by ewatrrr(24785)

 
Hi,
cola root beer sprite
under 21 years of age= 40 25 20 tot= 85
between 21 and 40= 35 20 30 tot = 85
over 40= 20 30 35 tot = 85
If one of the 255 subjects is randomly selected
P(over 40) = 85/255 = 1/3